OpenStudy (raheelafzaal):

If 2x + 2x + 2x = 12 then what is the value of x in 2x - 1 = 3?

OpenStudy (anonymous):

2x + 2x + 2x = 12 6x = 12 x = 12/6 x = 2 2x - 1 = 3 2(2) - 1 = 3 4 - 1 = 3 Is this what you were looking for?

OpenStudy (ikram002p):

is this a tricky question ?

OpenStudy (raheelafzaal):

If 2x + 2x + 2x = 12 then what is the value of x in 2x - 1 = 3? (a) 1 (b) 2 (c) 3 (d) 4